OpenStudy (anonymous):

If RQ  and  RS are midsegments of NOP , what can you conclude about QR and NP? Verify your results by finding x when QR = 3x + 2 and NP = 2x + 16.

OpenStudy (anonymous):

A. QR=NP; X=3 B. QR=2NP; X=9 C. QR= 1/2NP; X=3 D. QR= 1/3NP; X=9

OpenStudy (amoodarya):

\[\frac{QR}{NP}=\frac{1}{2}\]

OpenStudy (amoodarya):

\[\frac{3x+2}{2x+16}=\frac{1}{2} \\6x+4=2x+16 \\4x=12\\x=3\]

OpenStudy (amoodarya):

then 3x+2=9+2=11 2x+16=22 verified